Find the function that models the area (A) of an equillateral triange in terms of the length (X) of one of its sides:
Where am I going wrong...please help!
I've began solving as follows:
Length - X
height - h
Area of a triangle is = 1/2 (length times height)
Because this is an equilateral triange, the Length (X) is equal to its Height (H
Therefore,
A= 1/2 (X times X)
A= 1/2 x squared
so, the function is: A(X)=1/2x squared.
I know this is the wrong answer because my textbook gives the final answer but does a horrible job of explaining the solution. Please help! Thanks!

Answer by scott8148(6628)

"Because this is an equilateral triange, the Length (X) is equal to its Height (H"
___ this is NOT true

an altitude (height) of an equilateral triangle divides the triangle into two 30º-60º-90º triangles
___ with the length corresponding to the hypotenuse

the ratio of height to length is ___ sqrt(3) / 2

so ___ A = (1/2) (X) (X (sqrt(3)) / 2) = X^2 (sqrt(3)) / 4